OpenStudy (anonymous):

Which is the 23rd term of the arithmetic series 5, 14, 23, …

OpenStudy (anonymous):

do you know the formula? :)

OpenStudy (anonymous):

i know that it goes up by 9

Parth (parthkohli):

5 + (23 * 9) = 212 is the answer

OpenStudy (anonymous):

but she needs to know how to do it, just just the answer.

OpenStudy (anonymous):

cant be doesnt match up to the answers given

OpenStudy (anonymous):

yea kreshnik is right

Parth (parthkohli):

What are the choices?

OpenStudy (anonymous):

247 203 269 181

OpenStudy (anonymous):

i even counted them out and the closetest i got was 248

OpenStudy (anonymous):

\[\LARGE a_n=a_1+(n-1)d\] \[\LARGE d=a_n-a_{n-1 }\] so... we have: d=14-5=9 d=23-14=9 so d=9 and a_1=5 now we substitute: \[\LARGE a_{23}=\underbrace{5}_{a_1}+(23-1)\underbrace{9}_{d}\] so we have: \[\LARGE a_{23}=5+198\] can you do it now ? :)

OpenStudy (anonymous):

203

Parth (parthkohli):

I was about to say that haha

OpenStudy (anonymous):

lol ok. thanks.

OpenStudy (anonymous):

what if it gives me numbers other than a sequence like 25th term of the arithmetic sequence where a1 = 8 and a9 = 80

Parth (parthkohli):

@Kreshnik you genius!

OpenStudy (anonymous):

is it the same equation

OpenStudy (anonymous):

ahh.. Google crashed :) I hate this so much. You have no Idea. Here we write it again.

OpenStudy (anonymous):

what u mean write it again

OpenStudy (anonymous):

so we have: a1=8 and a9=80 we have to find a25 but we first have to find difference "d" how we find it? Easy ! let's solve for a9 \[\LARGE a_9=a_1+(9-1)d\] \[\LARGE 80=8+8d\] \[\LARGE 80-8=8d\] \[\LARGE 8d=72\] \[\LARGE d=\frac{72}{8}\] \[\LARGE d=9\] now we have difference d=9 we solve it for a25 \[\LARGE a_{25}=a_1+(25-1)d\] \[\LARGE a_{25}=8+(25-1)9\] and there you go :)

OpenStudy (anonymous):

what is it gives me the first and last term and asks for a term

OpenStudy (anonymous):

well sometimes you have a question like: a_3=20 a_8=50 find a_12 do you know how to do it ? :P